OpenStudy (anonymous):

a recipe for a pie calls for 1/3 cup of blackberries and 2/6 cup of raspberries .what is the total amount of berries needed for the recipe ? HELP HELP ??

OpenStudy (anonymous):

1/3+2/6=1/3+1/3=2/3 cup

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

yw